-8(6n + 2) - 6n = 2(-8n - 8) + n
Answer by ikleyn(52778)

.
-8(6n + 2) - 6n = 2(-8n - 8) + n
~~~~~~~~~~~~~~~~~~~~~~~~~~~

-48n - 16 - 6n = -16n - 16 + n,

-48n - 6n + 16n - n = -16 + 16,

-38n = 0,

n = 0.

Answer. n = 0.